OpenStudy (anonymous):

Rearrange to make x the subject. y = 2/3x^2

OpenStudy (anonymous):

GIVE THAT : y = 2/3(x^2) 3x^2y = 2 x^2 = 2/3y x = root of 2/3y

OpenStudy (anonymous):

i disagree y=2/3x^2 *multiply both sides by 3/2 3/2y=x^2 \[\sqrt{3/2y}=x\] \[x=\sqrt{3/2y}\]

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

kushashwa thought that x^2 is in the numerator

OpenStudy (anonymous):

hey dheeraj, do you write that as three over 2 times(secret x) y, or 3y over 2....or either?

OpenStudy (anonymous):

three over 2 times(secret x) y

OpenStudy (anonymous):

thanks dheeraj... exactly what I thought.

OpenStudy (anonymous):

hey give me a medal

OpenStudy (anonymous):

you got it wrong kushashwa...you mixed up the numerator and denominator. and dheeraj....it can be written as either!

OpenStudy (anonymous):

Oh sorry but u should understand it whether i had wrote wrong but the main thing is that the process was right and also i wrote that 's why y = 2/3(x^2) here i put x^2 in the bracket it was not my mistake but i accept it but 7181 u should right it clearly if u want a right answer in once hope u understand but i accept my mistake sorry for mistake :( :)